# Kang Chen♣,♡,∗, Zhihao Liu♢,♡,∗, Tonghe Zhang♯,⊳,∗, Zhen Guo, Si Xu, Hao Lin, Hongzhi Zang, & Xiang Li, Quanlu Zhang, Zhaofei Yu, Guoliang Fan, Tiejun Huang, Yu Wang♠,†, Chao Yu♠,♡,†

### v2

@@ Abstract
Vision-Language-Action (VLA) models enable robots to understand and perform complex tasks from multimodal input. Although recent work explores using reinforcement learning (RL) to automate the laborious data collection process in scaling supervised fine-tuning (SFT), applying large-scale RL to flow-based VLAs (e.g., $\pi_{0}$, $\pi_{0.5}$) remains challenging due to intractable action log-likelihoods from iterative denoising. We address this challenge with $\pi_{\texttt{RL}}$, an open-source framework for training flow-based VLAs in parallel simulation. $\pi_{\texttt{RL}}$ implements two RL algorithms: (1) Flow-Noise models the denoising process as a discrete-time MDP with a learnable noise network for exact log-likelihood computation. (2) Flow-SDE integrates denoising with agent-environment interaction, formulating a two-layer MDP that employs ODE-to-SDE conversion for efficient RL exploration. We evaluate $\pi_{\texttt{RL}}$ on LIBERO, ManiSkill, and MetaWorld benchmarks. On LIBERO, $\pi_{\texttt{RL}}$ boosts few-shot SFT models $\pi_{0}$ and $\pi_{0.5}$ from 57.6% to 97.6% and from 77.1% to 98.3%, respectively. On ManiSkill, we train $\pi_{\texttt{RL}}$ in 320 parallel environments, improving $\pi_{0}$ from 38.4% to 78.8% and $\pi_{0.5}$ from 40.1% to 90.8% across 4352 variations of pick-and-place task. On MetaWorld, RL is conducted over 50 different manipulation tasks and yields performance gains of 35.0% and 26.9% for $\pi_{0}$ and $\pi_{0.5}$ models, respectively. Overall, $\pi_{\texttt{RL}}$ achieves significant performance gains and stronger generalization over SFT-models, validating the effectiveness of online RL for flow-based VLAs.
Figure 1: Overview of $\pi_{\texttt{RL}}$. $\pi_{\texttt{RL}}$, an online RL framework featuring Flow-Noise and Flow-SDE two approaches, is designed to enhance the performance and generalization of SFT-aligned flow-based VLAs, represented by the $\pi_{0}$ and $\pi_{0.5}$. Experiments conducted on LIBERO, ManiSkill, and MetaWorld benchmarks demonstrate that $\pi_{\texttt{RL}}$ achieves significant gains over SFT models.
@@ 1 Introduction
Vision-Language-Action (VLA) models (din2025vision) have emerged as a leading solution for general-purpose robots, effectively bridging the gap between high-level multimodal reasoning and low-level physical control (firoozi2025foundation). Conditioned on sensor inputs and language commands, VLAs (team2024octo; kim2024openvla; black2024pi_0; intelligence2025pi05) can translate abstract instructions into executable robotic actions, thereby enabling intuitive and flexible human-robot interaction.
@@
The training methodology for VLAs follows the standard pre-training and supervised fine-tuning (SFT) paradigm as shown in Fig.˜1. Building on the pretrained Vision-Language Model (VLM) (touvron2023llama; beyer2024paligemma), VLAs are fine-tuned on large-scale, heterogeneous human demonstration datasets (o2024open; khazatsky2024droid), followed by SFT on the target task to align their capabilities with the specific embodiment and environment. However, reliance on SFT introduces a critical challenge: curating large-scale, high-quality expert trajectories is both laborious and costly (din2025vision), and the models obtained via SFT tend to overfit to expert demonstrations (liberoplus).
@@
Recent efforts (zang2025rlinf; li2025simplevla; tan2025riptvla; rl4vla) have explored expanding the VLA training process with reinforcement learning (RL), establishing a pre-training, SFT, and RL paradigm as shown in Fig.˜1, allowing VLAs to improve their performance beyond initial expert demonstrations through active environmental interaction and the development of more generalizable policies.
@@
However, these RL advances have been largely confined to autoregressive VLAs, featuring OpenVLA (kim2024openvla) and OpenVLA-OFT (kim2025openvlaoft), which employ discrete action decoders that generate output in an autoregressive or parallel fashion. This stands in stark contrast to diffusion- or flow-based VLAs, exemplified by the $\pi$ series models $\pi_{0}$ (black2024pi_0) and $\pi_{0.5}$ (intelligence2025pi05), which generate actions through iterative refinement in flow matching (lipman2022flow), offering the advantages of generating action chunks in high-frequency and performing highly dexterous tasks (black2024pi_0). Consequently, previous VLA-RL algorithms are incompatible with flow-based VLAs, and the fundamental challenge lies in how to characterize a logarithmic likelihood (hutchinson1989stochastic; chen2018neural) for the executed actions.
Contribution: We introduce $\pi_{\texttt{RL}}$, the first open-source framework for fine-tuning flow-based VLAs $\pi_{0}$ and $\pi_{0.5}$ with parallel online RL.
@@
To address the intractable log-likelihood estimation problem in flow matching, we propose two solutions: Flow-Noise and Flow-SDE. Flow-Noise integrates a learnable noise network into the denoising process and models this stage as a discrete-time Markov decision process (MDP) for exact log-likelihood estimation. Flow-SDE converts the ordinary differential equation (ODE) denoising process into a stochastic differential equation (SDE) while maintaining equivalent marginal distributions for exploration, and builds a two-layer MDP that couples the denoising process with policy-environment interaction, along with a hybrid ODE-SDE sampling technique for training acceleration. Given the formulated MDP and the exact log-likelihood computation, $\pi_{\texttt{RL}}$ undergoes further optimization via the proximal policy optimization (PPO) (schulman2017ppo) algorithm.
We conduct extensive experiments on the challenging multi-task benchmarks LIBERO (liu2023libero) and high-fidelity simulator ManiSkill (tao2024maniskill3) to evaluate the effectiveness of $\pi_{\texttt{RL}}$ optimization on the $\pi_{0}$ and $\pi_{0.5}$ models, with comprehensive findings summarized in Fig.˜1.
Results on LIBERO. $\pi_{\texttt{RL}}$ demonstrates substantial performance gains over the SFT baselines, with the average success rate of $\pi_{0}$ improving from 57.6% to 97.6%, and $\pi_{0.5}$ from 77.1% to 98.3%. Notably, on the LIBERO-Long task suite, $\pi_{\texttt{RL}}$ boosts the performance of the $\pi_{0.5}$ one-trajectory SFT model from 43.9% to 94.0%, surpassing the 92.4% performance of the all-trajectories SFT model. Moreover, we compare against group relative policy optimization (GRPO) (shao2024grpo) as an alternative policy gradient algorithm, with the comparison showing that PPO consistently outperforms GRPO across all task suites.
Results in ManiSkill. We train the policy to pick 16 types of objects and place them on 17 different receptacles in 16 photorealistic scenes, with a total of 4,352 combinations. $\pi_{\texttt{RL}}$ boosts the average success rate from 41.6% to 85.7% for $\pi_{0}$ and 41.9% to 84.8% for $\pi_{0.5}$, demonstrating $\pi_{\texttt{RL}}$’s ability to support large-scale multi-task RL. Additionally, we also conduct experiments on the SIMPLER benchmark (SIMPLER), the success rate was elevated from 67.2% to 86.7% for $\pi_{0}$ and from 59.24% to 79.1% for $\pi_{0.5}$.
Results in MetaWorld. Beyond the pick-and-place tasks detailed previously, we evaluate $\pi_{\texttt{RL}}$ on the MetaWorld MT50 benchmark to assess the capabilities of $\pi_{\texttt{RL}}$ across 50 diverse manipulation tasks of varying difficulty. With RL, the $\pi_{0}$ and $\pi_{0.5}$ models achieve success rates of 85.8% and 70.7%, exceeding the performance of the leading baseline SmolVLA (68.2%) (shukor2025smolvla).
@@
  • RL for flow-based VLAs. We introduce $\pi_{\texttt{RL}}$, the first online RL fine-tuning framework for flow-based $\pi$-series VLAs, featuring Flow-Noise and Flow-SDE, two distinct technical solutions that allow exact log-likelihood estimation in flow matching.
  • Superior Performance. We demonstrate significant performance improvements and enhanced generalization of $\pi_{\texttt{RL}}$ on the multi-task benchmarks LIBERO and ManiSkill.
  • Comprehensive Ablation. We conduct thorough ablation studies on RL algorithms, critic designs, noise injection strategies, MDP formulations, and hyperparameters within flow-based VLAs, providing empirical insights for future research on RL for flow-based VLAs.
  • Open-source Code and Models. We release all codes and model checkpoints to ensure reproducibility, hope thating that our study helps to advance further research in this field.
@@ 2.1 Vision-Language-Action Models
VLA models have recently achieved remarkable progress in robotics by integrating multimodal inputs to enable unified perception, reasoning, and control. This development has led to a series of architectures, including Octo (team2024octo), RT (brohan2022rt), OpenVLA, OpenVLA-OFT, $\pi_{0}$, $\pi_{0.5}$, and GR00T (bjorck2025gr00t). OpenVLA, which exemplifies the autoregressive VLA architecture, discretizes the action space into tokenized representations. This enables language-conditioned control by treating actions as part of the VLM’s vocabulary, but it inherently limits the resolution required for fine-grained motion. To achieve more dexterous and continuous physical behaviors, $\pi_{0}$ and $\pi_{0.5}$, as representatives of flow-based VLA architectures, introduce an action chunking architecture based on flow matching. This allows VLAs to model complex continuous action distributions, thereby achieving more dexterous physical behaviors.
In this work, we further fine-tune the $\pi$-series models with online RL algorithms, enhancing their performance and generalization capabilities through online interaction with the environment.
@@ 2.2 Online RL Fine-tuning for VLA Models
Recent research has increasingly focused on enhancing the performance and generalization of VLAs with online RL. For example, SimpleVLA-RL (li2025simplevla), building on the OpenVLA-OFT and GRPO, demonstrated that RL can improve long-horizon planning of VLA models under data scarcity. RL4VLA (rl4vla) empirically evaluated PPO, GRPO, and direct preference optimization (DPO) (rafailov2023dpo) with stage-based sparse rewards, finding PPO to yield superior performance. VLA-RL (lu2025vlarl) proposed a specialized robotic process reward model and enhanced the data processing pipeline. iRe-VLA (guo2025improving) proposed a framework that iterates between RL exploration and SFT updates. RIPT-VLA (tan2025riptvla) applied the REINFORCE leave-one-out (RLOO) (kool2018rloo) algorithm to the QueST (mete2024quest) and OpenVLA-OFT architectures. RLinf-VLA (yu2025rlinf; zang2025rlinf) provides a unified and efficient framework for scalable RL training of VLA models, supporting diverse VLA architectures such as OpenVLA and OpenVLA-OFT, multiple RL algorithms like PPO and GRPO, and various simulators including LIBERO and ManiSkill. These works demonstrate the effectiveness of RL fine-tuning VLA models.
While these approaches demonstrate the potential of applying online RL to VLAs, their application to flow-based VLAs is hindered by the challenge of exact log-likelihood estimation.
@@ 2.3 RL Fine-tuning for Flow Models
Integrating RL with flow models is a promising way to transcend the limitations of imitation learning. To this end, Flow-GRPO (liu2025flowgrpo) converts the deterministic ODE into an equivalent SDE to enable stochasticity exploration, a foundation upon which subsequent works like Mix-GRPO (li2025mixgrpo) and TempFlow-GRPO (he2025tempflow) further accelerate training through hybrid ODE-SDE rollouts. ReinFlow (zhang2025reinflow) injects learnable noise into the flow path and transforms it into a discrete-time Markov process with a tractable likelihood for stable policy gradient updates. Flow policy optimization (FPO) (mcallister2025fpo) reframes policy optimization as maximizing the advantage-weighted ratio of the conditional flow matching loss. Policy-Agnostic RL (PA-RL) parl effectively fine-tunes diverse diffusion and Transformer architectures by distilling critic-optimized actions into the policy via supervised learning. Diffusion steering via reinforcement learning (DSRL) (wagenmaker2025steering) refines the flow policy by performing RL in its latent-noise space, rather than modifying the policy parameters themselves.
While prior work has mostly focused on non-robotic tasks or small-scale, single-task robotics, we address the more challenging problem of fine-tuning large-scale flow-based VLAs for complex, multi-task robotic scenarios.
@@ 3.1 Problem Formulation
With the policy gradient surrogate (williams1992simple), the gradient of the return expectation can be approximated from sampled trajectories:
@@ 3.2 Flow-based Vision-Language-Action Model
Figure 2: Two optimization methods in $\pi_{\texttt{RL}}$. Flow-Noise adds learnable noise in a one-layer MDP (Fig.˜3), using the denoised joint likelihood for policy gradient. Flow-SDE builds a two-layer MDP with ODE-to-SDE conversion, and computes the likelihood directly.
@@ 4 Methodology
Existing VLA-RL approaches leverage base models such as OpenVLA for discrete actions and OpenVLA-OFT for continuous actions. To compute the action log-likelihood $\log\pi_{\theta}(a_{t}|s_{t})$, discrete models (rl4vla) apply softmax to the output logits, while continuous models (li2025simplevla) treat the action as a Gaussian distribution, employing a prediction head to estimate the variance. As for the flow-based VLAs, directly computing the exact likelihood (hutchinson1989stochastic) is inaccurate with few denoising steps. Moreover, the deterministic nature of its ODE sampling process precludes exploration, making its implementation within RL non-trivial. To this end, we propose Flow-Noise and Flow-SDE, two technical approaches that make flow-based VLAs amenable to RL, as depicted in Fig.˜2.
@@ 4.1 Flow-Noise
Inspired by Reinflow (zhang2025reinflow), we incorporate a learnable noise network into the flow matching denoising process and solve the problem within the standard one-layer MDP framework detailed in Sec.˜3.1. By modeling the denoising stage as a discrete MDP, we can directly compute the log-likelihood of the denoised sequence, enabling equivalent policy optimization via RL.
@@ 4.1.1 Stochasticity Injection
In Flow-Noise, we parameterize the noise schedule with a neural network, allowing the magnitude of the injected noise to be learned dynamically during training for greater flexibility, as shown in Fig.˜3. We focus on the generation process within a single environment timestep $t$. For notational simplicity, we omit the time subscript $t$, e.g., writing $\mathbf{A}^{\tau}$, and denote the predicted velocity $\mathbf{v}_{\theta}(\mathbf{A}^{\tau},\mathbf{o})$ as $\mathbf{v}^{\tau}$.
The step transition during the denoising process is modeled as an isotropic Gaussian distribution $p(\mathbf{A}^{\tau+\delta}|\mathbf{A}^{\tau})\sim\mathcal{N}(\mu_{\tau},\Sigma_{\tau})$, where the mean is determined by the forward Euler update of the original ODE and the variance is controlled by the learnable noise network $\theta^{\prime}$:
@@
Figure 3: Illustration for the noise injection on the flow matching, exemplified by $\pi_{0.5}$, which integrates image, language, and state information for unified VLM input.
@@ 4.1.2 Log-Likelihood Estimation
The primary challenge in applying policy gradient methods to flow-based VLAs stems from the intractable log-likelihood of the final executed action. In Flow-Noise, we address it by substituting the gradient of the joint log-likelihood of the entire denoising process into the policy optimization objective in Eq.˜2, which is theoretically grounded in Reinflow (zhang2025reinflow).
The inference process for action generation is discretized into $K$ uniform steps, which defines a sequence of time points $\{\tau_{0},\tau_{1},\dots,\tau_{K}\}$. With the step interval defined as $\delta=1/K$, the discrete timestep at the $k$-th point is $\tau_{k}=k\cdot\delta$, starting from $\tau_{0}=0$ and culminating at $\tau_{K}=1$. Given the observation $\mathbf{o}$, the exact and tractable log probability for the entire denoising sequence $\mathcal{A}=(\mathbf{A}^{0},\dots,\mathbf{A}^{1})$ is depicted in Fig.˜2 and formulated as:

### v3

# Kang Chen2,6,∗    Zhihao Liu3,6,∗    Tonghe Zhang4,∗,♯    Zhen Guo5    Si Xu5    Hao Lin5    Hongzhi Zang1    Xiang Li5    Bingwen Wei1    Jiakai Zhou1    Quanlu Zhang5    Zhaofei Yu2    Guoliang Fan3    Tiejun Huang2    Yu Wang    Chao Yu

@@ Abstract
Vision-Language-Action (VLA) models enable robots to understand and perform complex tasks from multimodal input. Although recent work explores using reinforcement learning (RL) to automate the laborious data collection process in scaling supervised fine-tuning (SFT), applying RL to large-scale flow-based VLAs (e.g., $\pi_{0}$, $\pi_{0.5}$) remains challenging due to intractable action log-likelihoods raised from flow matching. We address this challenge with $\pi_{\texttt{RL}}$, featuring two technical approaches: (1) Flow-Noise models the denoising process as a discrete-time MDP with a learnable noise network for exact log-likelihood computation. (2) Flow-SDE integrates denoising with agent-environment interaction, formulating a two-layer MDP that employs ODE-to-SDE conversion for efficient RL exploration. We evaluate $\pi_{\texttt{RL}}$ across various benchmarks, with experiments demonstrating that RL yields significant performance improvements in both in-distribution and out-of-distribution settings.
Machine Learning, ICML
Figure 1: $\pi_{\texttt{RL}}$: An online RL framework for flow-based VLAs. Incorporating two solutions, Flow-Noise and Flow-SDE, $\pi_{\texttt{RL}}$ enhances the performance and generalization of SFT-aligned models across extensive ID benchmarks and OOD settings. Refined with RL, few-shot SFT policies achieve performance comparable to full dataset baselines. Additionally, we facilitate seamless zero-shot sim-to-real transfer by constructing a simulator with 3D Gaussian Splatting as the rendering engine to narrow the visual domain gap.
@@ 1 Introduction
Vision-Language-Action (VLA) models (Din et al., 2025) have emerged as a leading solution for general-purpose robots, effectively bridging the gap between high-level multimodal reasoning and low-level physical control (Firoozi et al., 2025). Conditioned on sensor inputs and language commands, VLAs (Team et al., 2024; Kim et al., 2024; Black et al., 2024; Intelligence et al., 2025) can translate abstract instructions into executable robotic actions, thereby enabling intuitive and flexible human-robot interaction.
@@
The training methodology for VLAs follows the standard pre-training and supervised fine-tuning (SFT) paradigm as shown in Fig.˜1. Building on the pretrained Vision-Language Model (VLM) (Touvron et al., 2023; Beyer et al., 2024), VLAs are fine-tuned on large-scale, heterogeneous human demonstration datasets (O’Neill et al., 2024; Khazatsky et al., 2024), followed by SFT on the target task to align their capabilities with the specific embodiment and environment. However, reliance on SFT introduces a critical challenge: curating large-scale, high-quality expert trajectories is both laborious and costly (Din et al., 2025). Besides, models obtained via SFT tend to overfit to expert demonstrations (Fei et al., 2025), with their performance fundamentally constrained by the quality of expert demonstrations.
@@
Recent efforts (Zang et al., 2025; Li et al., 2025a; Tan et al., 2025; Liu et al., 2025b) have explored expanding the VLA training process with reinforcement learning (RL), establishing a pre-training, SFT, and RL paradigm as shown in Fig.˜1, allowing VLAs to improve their performance beyond expert demonstrations through environmental interaction and develop more generalizable policies.
@@
However, these RL advances have been largely confined to autoregressive VLAs, featuring OpenVLA (Kim et al., 2024) and OpenVLA-OFT (Kim et al., 2025), which employ discrete action decoders that generate output in an autoregressive or parallel fashion. This stands in stark contrast to flow-based VLAs, exemplified by the $\pi$ series models, which generate actions through iterative refinement in flow matching (Lipman et al., 2022), offering the advantages of generating action chunks in high-frequency and performing highly dexterous tasks (Black et al., 2024). Consequently, previous VLA-RL algorithms are incompatible with flow-based VLAs, and the fundamental challenge lies in how to characterize a logarithmic likelihood (Hutchinson, 1989; Chen et al., 2018) for the executed actions.
@@
In this paper, we introduce $\pi_{\texttt{RL}}$, a framework designed for fine-tuning flow-based VLAs with online RL algorithms. To address the intractable log-likelihood estimation problem in flow matching, we propose two solutions. Flow-Noise integrates a learnable noise network into the denoising process and models this stage as a discrete-time Markov decision process (MDP) for exact log-likelihood estimation. Flow-SDE converts the ordinary differential equation (ODE) denoising process into a stochastic differential equation (SDE) while maintaining equivalent marginal distributions for exploration, and builds a two-layer MDP that couples the denoising process with policy-environment interaction. Given the formulated MDP and the exact log-likelihood computation, $\pi_{\texttt{RL}}$ undergoes further optimization via the proximal policy optimization (PPO) (Schulman et al., 2017).
We conduct extensive experiments on various benchmarks to evaluate the effectiveness of $\pi_{\texttt{RL}}$ on $\pi_{0}$ (Black et al., 2024) and $\pi_{0.5}$ (Intelligence et al., 2025) models. Across all benchmarks, the proposed framework consistently yields substantial performance gains over SFT baselines. Furthermore, out of distribution evaluations confirm that our model yields genuine policy enhancement rather than narrow overfitting on the target environment.
@@
  • RL for flow-based VLAs. We introduce $\pi_{\texttt{RL}}$, an online RL fine-tuning framework with Flow-Noise and Flow-SDE formulations for flow-based VLAs.
  • Superior Performance. We demonstrate significant performance improvements and enhanced generalization of $\pi_{\texttt{RL}}$ across various benchmarks.
  • Comprehensive Ablation. We conduct thorough ablation studies, offering empirical insights to guide future RL research on flow-based VLAs.
  • Open-source Code and Models. We release all codes to ensure reproducibility, hoping that our study helps to advance further research in this field.
@@ 2.1 Vision-Language-Action Models
VLA models have recently achieved remarkable progress in robotics by integrating multimodal inputs to enable unified perception, reasoning, and control. This development has led to a series of architectures, including Octo (Team et al., 2024), RT (Brohan et al., 2022), OpenVLA, OpenVLA-OFT, $\pi_{0}$, $\pi_{0.5}$, and GR00T (Bjorck et al., 2025).
@@ 2.2 Online RL Fine-tuning for VLA Models
Recent research has increasingly focused on enhancing the performance and generalization of VLAs with online RL. For example, SimpleVLA-RL (Li et al., 2025a), building on the OpenVLA-OFT and GRPO, demonstrated that RL can improve long-horizon planning of VLA models under data scarcity. RL4VLA (Liu et al., 2025b) empirically evaluated PPO, GRPO, and direct preference optimization (DPO) (Rafailov et al., 2023) with stage-based sparse rewards. RLinf-VLA (Yu et al., 2025; Zang et al., 2025) provides a unified and efficient framework for scalable RL training of VLA models. These works demonstrate the effectiveness of RL fine-tuning VLA models.
@@ 2.3 RL Fine-tuning for Flow Models
Integrating RL with flow models is a promising way to transcend the limitations of imitation learning. To this end, Flow-GRPO (Liu et al., 2025a) converts the deterministic ODE into an equivalent SDE to enable stochasticity exploration, a foundation upon which subsequent works like Mix-GRPO (Li et al., 2025b) and TempFlow-GRPO (He et al., 2025) further accelerate training through hybrid ODE-SDE rollouts. ReinFlow (Zhang et al., 2025) injects learnable noise into the flow path and transforms it into a discrete-time Markov process with a tractable likelihood for stable policy gradient updates. Flow policy optimization (FPO) (McAllister et al., 2025) reframes policy optimization as maximizing the advantage-weighted ratio of the conditional flow matching loss.
@@ 3.1 Problem Formulation
With the policy gradient surrogate (Williams, 1992), the gradient of the return expectation can be approximated from sampled trajectories:
@@ 4 Methodology
Existing VLA-RL approaches leverage base models such as OpenVLA for discrete actions and OpenVLA-OFT for continuous actions. To compute the action log-likelihood $\log\pi_{\theta}(a_{t}|s_{t})$, discrete models (Liu et al., 2025b) apply softmax to the output logits, while continuous models (Li et al., 2025a) treat the action as a Gaussian distribution, employing a prediction head to estimate the variance. As for the flow-based VLAs, directly computing the exact likelihood (Hutchinson, 1989) is inaccurate with few denoising steps. Moreover, the deterministic nature of its ODE sampling process precludes exploration, making its implementation within RL non-trivial. To this end, we propose Flow-Noise and Flow-SDE, two technical approaches that make flow-based VLAs amenable to RL.
@@ 4.1 Flow-Noise
Inspired by Reinflow (Zhang et al., 2025), we incorporate a learnable noise network into the flow matching denoising process and solve the problem within the standard one-layer MDP framework detailed in Sec.˜3.1. By modeling the denoising stage as a discrete MDP, we can directly compute the log-likelihood of the denoised sequence, enabling equivalent policy optimization via RL.
@@ 4.1.1 Stochasticity Injection
In Flow-Noise, we parameterize the noise schedule with a neural network, allowing the magnitude of the injected noise to be learned dynamically during training for greater flexibility, as shown in Fig.˜2. We focus on the generation process within a single environment timestep $t$. For notational simplicity, we omit the time subscript $t$, e.g., writing $\mathbf{A}^{\tau}$, and denote the predicted velocity $\mathbf{v}_{\theta}(\mathbf{A}^{\tau},\mathbf{o})$ as $\mathbf{v}^{\tau}$.
The step transition during the denoising process is modeled as a Gaussian distribution $p(\mathbf{A}^{\tau+\delta}|\mathbf{A}^{\tau})\sim\mathcal{N}(\mu_{\tau},\Sigma_{\tau})$, where the mean is determined by the forward Euler update of the original ODE and the variance is controlled by the learnable noise network $\theta^{\prime}$:
@@
Figure 2: Illustration for the noise injection in $\pi_{\texttt{RL}}$.
@@ 4.1.2 Log-Likelihood Estimation
The primary challenge in applying policy gradient methods to flow-based VLAs stems from the intractable log-likelihood of the final executed action. In Flow-Noise, we address it by substituting the gradient of the joint log-likelihood of the entire denoising process into the policy optimization objective in Eq.˜2, which is theoretically grounded in Reinflow (Zhang et al., 2025).
The inference process for action generation is discretized into $K$ uniform steps, which defines a sequence of time points $\{\tau_{0},\tau_{1},\dots,\tau_{K}\}$. With the step interval defined as $\delta=1/K$, the discrete timestep at the $k$-th point is $\tau_{k}=k\cdot\delta$, starting from $\tau_{0}=0$ and culminating at $\tau_{K}=1$. Given the observation $\mathbf{o}$, the exact and tractable log probability for the entire denoising sequence $\mathcal{A}=(\mathbf{A}^{0},\dots,\mathbf{A}^{1})$ is depicted in Fig.˜3 and formulated as: